Oolie Corp. paid $6 million to acquire a 25% interest in Ketchum Corp. common stock on January 1, 2014. This investment results in Oolie having significant influence over Ketchum’s operations. The book values of Ketchum’s reported net assets approximate their fair values, except for $2 million in identifiable unrecorded intangible assets that have a four-year remaining useful life on January 1, 2014. Ketchum has no unidentifiable intangibles (i.e., there is no implied goodwill). During 2014, Ketchum reports net income of $900,000. During 2014, Ketchum also reports, as part of other comprehensive income, $30,000 in unrealized losses on available for sale (AFS) debt securities that were purchased in 2014. Ketchum declared and paid dividends of $240,000 during 2014. (Hint: you might want to check out in FASB ASC sections 323-10-35-18 and 323-10-45-3.)

  1. Prepare Oolie’s journal entries for its investment in Ketchum for the year ended December 31, 2014.
  2. For the year ended December 31, 2014, what is the effect on the net income of Oolie of equity method accounting for the investment in Ketchum?
  3. For the year ended December 31, 2014, what is the effect on the comprehensive income of Oolie of equity method accounting for the investment in Ketchum?
Answer #1
in $Mn
Date Account Debit Credit
01-01-2014 Investment $ 6.00
Bank $ 6.00
Investment purchased in Ketchum Corp.
31-12-2014 Investment $ 0.23
Income $ 0.23
Share of oolie recorded in income of Ketchum Corp.
$0.9 * 0.25 = $0.23
31-12-2014 Depreciation (Income) $ 0.13
Investment $ 0.13
Depreciation on intangible asset recorded)
$2 /4 years * 0.25 = $0.13 mn
31-12-2014 Dividend(Income) $ 0.06
Investment $ 0.06
Share of dividend recorded
$0.24*0.25 = $0.06mn
Investment Value as on 31/12/14
Purchase price $ 6.00
Add: Net income share in Ketchum Corp $ 0.23
Less: Depreciation on intangible asset $ 0.13
Less: Dividend $ 0.06
Less: Unrealized loss $ 0.08
$ 5.97
Proportionate share of Net income of 2014 of Ketchum Corp. will be added in the investment value of Oolie corp.
$0.9mn * 0.25 = $0.23 mn will be added in incoem of Oolie
Unrealized gain or loss of Ketchum Corp. will also be recorded in books of Oolie Corp.
Loss of $30000 will also be recorded
$30000 * 25% = $7500